# Materials Map Integrating Experimental and Computational Data through Graph-Based Machine Learning for Enhanced Materials Discovery

### v1

@@ I Introduction
Materials informatics has emerged as a transformative approach to accelerate materials development by integrating data science with materials science [1, 2, 3, 4, 5]. Through the application of artificial intelligence (AI) and machine learning (ML), researchers can efficiently identify and design materials with desired properties, significantly reducing the time and cost associated with traditional experimental methods. The success of such approaches relies heavily on the availability of high-quality, large-scale datasets. Initiatives such as the Materials Project [6] and AFLOW [7] have been instrumental in systematically collecting and organizing results from first-principles calculations conducted globally. These databases, when combined with advanced data science techniques, enable the development of machine learning models capable of predicting material properties with remarkable accuracy [8]. A critical challenge persists in bridging the gap between theoretical predictions and practical applications. While computational databases provide extensive datasets derived from first-principles calculations, experimental data remains sparse, inconsistent, and often lacks the structural information necessary for advanced modeling. This disparity between computational and experimental data presents a significant obstacle to the full realization of materials informatics potential.
@@
In our previous study [25], we proposed a novel framework to integrate experimental and computational datasets to create a comprehensive data for materials informatics. As briefly described in the SI, our approach consisted of the following steps: preprocessing experimental data, training a machine learning model, and applying the trained model to predict experimental data for the compositions registered in the computational data. As a result, we obtained a dataset of the predicted experimental values of material properties () with structural information of the corresponding materials.
@@ II Method
As schematically described in Figure 1, we here construct materials maps by MDL, which implements graph-based prepresentation of materials properties, deep learning, and dimensional reduction by the t-SNE algorithm [8]. We use the dataset extracted by our previous study, described in Ref. [25] and briefly summarized in SI. This dataset comprises over 1,000 materials, each represented by its structural information and predicted experimental values. The structural information includes atomic positions and lattice parameters, enabling the construction of graph-based representation of materials by MDL [8], which provides a flexible environment for material property prediction. As schematically described in Figure 1, the model defined in MDL is composed of input, embedding, graph convolution, pooling, dense, and output layers. The entire model is trained using material structures as input and the corresponding predicted values as output. The input layer primarily extracts basic structural information, such as atomic positions, types, and bond distances, using the Atomic Simulation Environment (ASE) framework. The obtained data serves as the foundation for constructing graph structures and developing advanced machine learning models, including deep learning approaches. MDL supports the construction of material maps using t-SNE with the data from the first dense layer, which is placed after GC layer as schematically described in Figure 2. This configuration allows the model to capture both local and long-range structural correlations learned by the GC layer.
@@ III.2 Contribution of GC Layer in MPNN Architecture for Map Construction
The framework of DeeperGATGNN allows machine learning modeling with large N\̃\ citeOmee2022-tk. With increasing N, prediction accuracy and structural feature learning are enhanced, while memory usage for computation increases dramatically especially when large datasets are used for the analysis\̃\ citeOmee2022-tk. Therefore, the use of MDL with MPNN architecture may be limited to cases where the dataset size is relatively small or when computing resources with sufficient memory size are available.
@@ III.3 Model Dependency of Material Maps
Similar discussions are found in Ref. [8], which is the first report of MDL, but the differences in maps obtained by different models were not significant compared to the present case. We attribute the differences in these model dependencies in map construction to data filtering. Both studies use the same database, Materials Project, while, in this study, we applied a filter based on material stability, band gap, and atomic elements. This resulted in the selection of similar materials, which facilitated feature extraction by machine learning models, leading to material maps with clearer structures.
@@ IV.1 Relationship between material property prediction and material map construction
The machine learning model defined in MDL predicts materials properties and generates material maps as part of a continuous process. (Note that this prediction by MDL is different from the prediction of the experimental values, described in Ref. [25].) The creation of high-quality material maps indicates the effective extraction of materials features, and thus leads us to anticipate enhanced accuracy in material property predictions. This logic motivate us to evaluate the model performance for the material property prediction by three standard metrics: the coefficient of determination (R²), mean squared error (MSE), and mean absolute error (MAE) with the data divided into training (80), testing (15), and validation (5) subsets. The evaluation results with the validation dataset, summarized in Table 1, leads that, while the MPNN produces the most structurally interpretable maps, its prediction accuracy (R2 = 0.610) is the lowest among all the models. This unexpected result leads us to conclude that the clear structural organization displayed in the MPNN-derived maps does not necessarily translate to higher predictive performance of the materials properties.
@@ IV.2 Improvement of the interpretability of material maps
To investigate the distribution of materials in the map, we applied -means clustering analysis to the map with =10 clusters as shown in Figure 6(f). The label for each cluster was assigned by sorting based on the average of predicted values, allowing us to identify regions of high-performance materials. To understand the materials included in each cluster, we counted the number of elements in the compositions of the materials included in each cluster. The elements with less than 10 of the compositional ratio was excluded to account for the elements that make a major contribution to the material properties. The results for each cluster are plotted using the Periodic Trend Plotter (https://github.com/Andrew-S-Rosen/periodic_trends) in Figure 7. In order to find the characteristics of the two branches, labeled as BR1 and BR2 in Figure 2, we compare the elemental distribution for clusters 9 and 10, which are placed at the edge of each branch. We find high similarities between these two plots, such as the high frequencies of Te atom, supporting an improved understanding of the obtained material maps.
@@ IV.4 Prospects for future studies
Creating materials maps require substantial computational resources, especially for tasks such as dimensionality reduction by the t-SNE architecture and the deep-learning modeling defined in MDL. In this study, we employed the model and the hyperparameters given in MDL due to computational constraints. For wider and better applicability, fine-tuning of the model may be important. Recent advancements in high-performance computing (HPC) and cloud-based GPU/TPU clusters will help mitigate these computational challenges, enabling broader adoption of such methods.
@@ V Conclusion
In this study, we developed and evaluated a novel approach for materials exploration by constructing a materials map that combines graph theory-based model and deep learning. A key feature of our work is the use of materials data from our previous research, which incorporated experimentally predicted values of material properties () along with structural information. The materials maps, generated using the MatDeepLearn (MDL) framework, were evaluated through distribution analysis of of nearest-neighbor distances between data points by kernel density estimation. Results showed that among the graph theory-based models supported by MDL, the Message Passing Neural Network (MPNN) provided the most efficient structuring of material properties. Additionally, a systematic evaluation controlled training by the graph convolution layers of MDL demonstrated their significant contribution to the organization of the materials map.
Acknowledgments
@@
VI Supplemental information
VI.1 Data Preparation and Machine Learning
VI.1.1 Cleaning of Data in StarryData2
@@
VI.1.2 Machine Learning Modeling of Experimental Data
@@
VI.1.3 Machine Learning Modeling of Calculation Data
@@
VI.2 Correlation analysis of parameters
@@
VI.3 Dimensionality Reduction for Materials Map Construction
@@
VI.4 Hardware and Software Environment

### v2

@@ I Introduction
Materials informatics has emerged as a transformative approach to accelerate materials development by integrating data science with materials science [1, 2, 3, 4, 5]. Through the application of artificial intelligence (AI) and machine learning (ML), researchers can efficiently identify and design materials with desired properties, significantly reducing the time and cost associated with traditional experimental methods. The success of such approaches relies heavily on the availability of high-quality, large-scale datasets. Initiatives such as the Materials Project (MP) [6] and AFLOW [7] have been instrumental in systematically collecting and organizing results from first-principles calculations conducted globally. These databases, when combined with advanced data science techniques, enable the development of machine learning models capable of predicting material properties with remarkable accuracy [8]. A critical challenge persists in bridging the gap between theoretical predictions and practical applications. While computational databases provide extensive datasets derived from first-principles calculations, experimental data remains sparse, inconsistent, and often lacks the structural information necessary for advanced modeling. This disparity between computational and experimental data presents a significant obstacle to the full realization of materials informatics potential.
@@
In our previous study [25], we proposed a novel framework to integrate experimental and computational datasets to create a comprehensive data for materials informatics. As briefly described in SI, our approach consisted of the following steps: preprocessing experimental data, training a machine learning model, and applying the trained model to predict experimental data for the compositions registered in the computational data. As a result, we obtained a dataset of the predicted experimental values of material properties () with structural information of the corresponding materials.
@@ II Method
As schematically described in Figure 1, we here construct materials maps by MDL, which implements graph-based resentation of materials properties, deep learning, and dimensional reduction by the t-SNE algorithm [8]. We use the dataset extracted by our previous study, described in Ref. [25] and briefly summarized in SI. This dataset comprises over 1,000 materials, each represented by its structural information and predicted experimental values. The structural information includes atomic positions and lattice parameters, enabling the construction of graph-based representation of materials by MDL [8], which provides a flexible environment for material property prediction. As schematically described in Figure 1, the model defined in MDL is composed of input, embedding, graph convolution, pooling, dense, and output layers. The entire model is trained using material structures as input and the corresponding predicted values as output. The input layer primarily extracts basic structural information, such as atomic positions, types, and bond distances, using the Atomic Simulation Environment (ASE) framework. The obtained data serves as the foundation for constructing graph structures and developing advanced machine learning models, including deep learning approaches. MDL supports the construction of material maps using t-SNE with the data from the first dense layer, which is placed after GC layer as schematically described in Figure 2. This configuration allows the model to capture both local and long-range structural correlations learned by the GC layer.
@@ III.2 Contribution of GC Layer in MPNN Architecture for Map Construction
The framework of DeeperGATGNN allows machine learning modeling with large N [15]. With increasing N, prediction accuracy and structural feature learning are enhanced, while memory usage for computation increases dramatically especially when large datasets are used for the analysis [15]. Therefore, the use of MDL with MPNN architecture may be limited to cases where the dataset size is relatively small or when computing resources with sufficient memory size are available.
@@ III.3 Model Dependency of Material Maps
Similar discussions are found in Ref. [8], which is the first report of MDL. This study also compares material maps generated with different models; however, the differences in the obtained maps were not as significant compared to the present case. We attribute the difference in these two studies to the data screening applied in our study. Both studies use the same database, Materials Project, while, in this study, we applied a filter based on material stability, band gap, and atomic elements. This screening results in the selection of similar materials, which facilitated feature extraction by machine learning models, leading to material maps with clearer structures.
@@ IV.1 Relationship between material property prediction and material map construction
The machine learning model defined in MDL predicts materials properties and generates material maps as part of a continuous process. (Note that this prediction by MDL is different from the prediction of the experimental values, described in Ref. [25].) The creation of high-quality material maps indicates the effective extraction of materials features, and thus leads us to anticipate enhanced accuracy in material property predictions. This logic motivate us to evaluate the model performance for the material property prediction by three standard metrics: the coefficient of determination (R2), mean squared error (MSE), and mean absolute error (MAE) with the data divided into training (80), testing (15), and validation (5) subsets. The evaluation results with the validation dataset, summarized in Table 1, leads that, while the MPNN produces the most structurally interpretable maps, its prediction accuracy (R2 = 0.610) is the lowest among all the models. This unexpected result leads us to conclude that the clear structural organization displayed in the MPNN-derived maps does not necessarily translate to higher predictive performance of the materials properties.
@@ IV.2 Improvement of the interpretability of material maps
To investigate the distribution of materials in the map, we applied -means clustering analysis to the map with =10 clusters as shown in Figure 6(f). The label for each cluster was assigned by sorting based on the average of predicted values, allowing us to identify regions of high-performance materials. To understand the materials included in each cluster, we counted the number of elements in the compositions of the materials included in each cluster. The elements with less than 10 of the compositional ratio was excluded to account for the elements that make a major contribution to the material properties. The results for each cluster are plotted using the Periodic Trend Plotter (https://github.com/Andrew-S-Rosen/periodic_trends) in Figure 7. In each cluster, we find distinct characteristic distributions of S, Se, and Te atoms. Clusters containing a large number of materials with low zT values (1–7) show a high frequency of S and Se occurrences. On the other hand, clusters containing a large number of materials with high zT values (8–10) exhibit a higher frequency of Te occurrences. In order to find the characteristics of the two branches, labeled as BR1 and BR2 in Figure 2, we compare the elemental distribution for clusters 9 and 10, which are placed at the edge of each branch. We find high similarities between these two plots, such as the high frequencies of Te , Sb, Ti, and Ag atoms. The detailed discussion in terms of the atomic distributions supports our improved understanding of the obtained material maps.
@@ IV.4 Prospects for future studies
Creation of materials maps requires substantial computational resources, especially for tasks such as dimensionality reduction by the t-SNE architecture and the deep-learning modeling defined in MDL. In this study, we employed the model and the hyperparameters given in MDL due to computational constraints. For wider and better applicability, fine-tuning of the model may be important. Recent advancements in high-performance computing (HPC) and cloud-based GPU/TPU clusters will help mitigate these computational challenges, enabling broader adoption of such methods.
@@ V Conclusion
In this study, we developed and evaluated a novel approach for materials exploration by constructing a materials map that combines graph theory-based model and deep learning. A key feature of our work is the use of materials data from our previous research, which incorporated experimentally predicted values of material properties () along with structural information. The materials maps, generated using the MatDeepLearn (MDL) framework, were evaluated through distribution analysis of nearest-neighbor distances between data points by kernel density estimation. Results showed that among the graph theory-based models supported by MDL, the Message Passing Neural Network (MPNN) provided the most efficient structuring of material properties. Additionally, a systematic evaluation controlled training by the graph convolution layers of MDL demonstrated their significant contribution to the organization of the materials map.
VI Acknowledgments
@@
VII Supplemental information
VII.1 Data Preparation and Machine Learning
VII.1.1 Cleaning of Data in StarryData2
@@
VII.1.2 Machine Learning Modeling of Experimental Data
@@
VII.1.3 Machine Learning Modeling of Calculation Data
@@
VII.2 Correlation analysis of parameters
@@
VII.3 Dimensionality Reduction for Materials Map Construction
@@
VII.4 Hardware and Software Environment